Calculate the next limit, if it doesn’t exist, then prove it.
2 y (b) lim (x,y)→(0,0) sin y + ln(1 + x2)
Answer #1

lim 0+ y - siny 2sinycosy ,训cos2y = lim put y-mx x2 + m2x2 lim = lim 2x(1+m 2) 2(1+m (1+x2)2-2x(2x) 2m2 cos(2mx)+ along y=mx2